The subsidiary of a U.S. corporation located in Country Y generated income of $1,000,000 on which...

  1. The subsidiary of a U.S. corporation located in Country Y generated income of $1,000,000 on which it paid $200,000 in taxes to Country Y. The subsidiary paid a dividend to the U.S. parent of $150,000. How much tax is currently owed to the U.S. government if the federal tax rate is 35%? Would the response be different if we were working with a branch instead of a subsidiary?
Answer #1

Answer is 28125

Income of Subsidiary 1000000
Less: Taxes Paid outside 200000
Earnings available for distribution   800000
Declared as dividend

150000

Dividends received creditable

Deemed-paid credit =

= 1500000 / 800000 x 200000

= 37500

Now, Gross amount deemed to be recieved in US = Amount Paid as dividend + proportionate amount of tax credit

= 150000 + 37500 = 187500

Us Taxes computed on gross amount received in US = 187500 x 35 % ( US Tax rate)

= 65625

Now Net amount payable in US Taxes = Amount of Tax in US - Deemed Paid credit

= 65625 - 37500 , = 28125